phone call arrive at the rate of 48 per our at the reservation desk for sri lankan airways.Find the probability of receiving three call in a five minute interval of time


1
Expert's answer
2021-02-04T14:03:15-0500

Poisson distribution:

"P(X=x) = \frac{e^{-λ}λ^x}{x!} \\

λ = \frac{48}{60} = 0.8 \; calls \; per \;min \\

λ = 0.8 \times 5 = 4 \;calls \;per 5 \; min \\

P(X=3) = \frac{e^{-4} \times 4^3}{3!} = 0.1954"

Answer: 0.1954